I have 500 dollars in the bank, I will continue to put 100 dollars into the bank account monthly. I have a compound interest of .8%. What is the equation to find the solution at any given time? What is the total amount after 10 years?
Answer by DrBeeee(684)

The formula is
(1) B = P*(1+r)^m
Where B = your balance of money ($) at the end of m months when P is the amount you start with ($500 in your case) and r is the MONTHLY interest rate expressed as a decimal. Apply (1) to your case we get
(2) B = 500*(1.008)^m
After 10 years or 120 months you will have
(3) B = 500*(1.008)^120 or
(4) B = 500*(2.6017398) or
(5) B = 1300.8699 or
(6) B = $1300.87
If it was not compounded monthly you would get
(7) b = 500*(1.096)^10 or
(8) b = 500(2.5009531) or
(9) b = $1250.48
Only about $50 less. The compounding is over-rated don't you think?